# Reshaping Perception Through Technology: From Ancient Script to Large Language Models

### v1

@@ Abstract
Large language models are reshaping how we create and access information, yet we typically view perception as merely reactive to stimuli, overlooking how the physical qualities of different media uniquely shape cognition. Drawing on Marshall McLuhan’s insight that “the medium is the massage,” we trace a lineage of technologies – from DNA and the nervous system to language, writing, music, and now LLMs – that mold perception in distinct ways. We observe that as technologies become more advanced and decoupled from our physiology, they introduce both greater creative potential and greater risk: they enable more efficient play, storage, and transmission, while also introducing artificiality and the potential for inauthenticity and manipulation. This tension is particularly acute with LLMs, which allow rapid, playful generation of content increasingly indistinguishable from human-created work. Noting that humans have a recurring tendency to project intelligence onto novel technologies (a pattern visible in ancient responses to writing), we argue that AI should be framed not as a competitor but as a medium that reshapes perceptual skills and enables new forms of creativity.
@@ 3  Folding and Unfolding in the Brain
Given this understanding, brain dynamics more closely mirror the embryologic logic of differentiation and flexible unfolding described above than the operation of an electrical circuit. Classical research in perceptual decision-making reveals that neural activity accumulates over time as organisms prepare to act. This accumulation is driven by motivation but refined through inhibitory signals from neurons associated with competing or irrelevant decision options. Critically, the spontaneous variability inherent in this process affords the potential for the exploration of novel patterns, and through neuroplasticity, the potential to write and rewrite the synaptic structures that channel neuro-electrical activity in the first place. Seen in this light, our neural network is a “written artifact”, representing the history of useful sculptings of an endless ocean of spontaneous activity – the manifestation of evolutionary, developmental, and experiential unfolding. This written artifact is similar to what German philosopher Friedrich Schelling called architecture as “frozen music” – in our example, the music is the ongoing spontaneous activity itself and the synaptic strengths the resultant architecture (von Schelling, 1845).
@@ 4  The Medium Shapes the Message
This discussion highlights our emphasis on the word medium: the defining mediation between subject and object. A medium’s spatiotemporal features are key in this mediation process. Canadian philosopher Marshall McLuhan famously coined the phrase, “The medium is the message”, observing that the unique properties associated with a novel medium profoundly shape our conscious experience – they literally become stretched “extensions” of our perceptual process, as he put it. Consequently, the real message is not a code in a channel of information, but rather how a medium’s unique characteristics directly shape our experience from the ground up. For this reason, McLuhan later adopted the title “the medium is the massage”, to imply how the unique features of mediums “massage” our conscious experience (McLuhan and Fiore, 1967).

### v2

@@ Abstract
As large language models reshape how we create and access information, questions arise about how to frame their role in human creative and cognitive life. We argue that AI is best understood not as artificial intelligence but as a new medium—one that, like writing before it, reshapes perception and enables novel forms of creativity. Drawing on Marshall McLuhan’s insight that “the medium is the massage,” we trace a lineage of technologies—from DNA and the nervous system to symbols, writing, and now LLMs—that mold cognition through a shared logic of flexible unfolding and co-creation. We observe that as technologies become more externalized and decoupled from physiology, they introduce both greater creative potential and greater risk of inauthenticity and manipulation. This tension is acute with LLMs, but not unprecedented: ancient responses to writing reveal a recurring human tendency to project intelligence onto powerful new media. Rather than viewing AI as a competitor, we propose framing it as a medium that foregrounds artistic skills: aesthetic judgment, curation, and the articulation of vision. We discuss implications for education, creative practice, and how society might adapt to this new medium as it did to writing.
@@ 3  Folding and Unfolding in the Brain
Given this understanding, brain dynamics more closely mirror the embryologic logic of differentiation and flexible unfolding described above than the operation of an electrical circuit. Classical research in perceptual decision-making reveals that neural activity accumulates over time as organisms prepare to act. This accumulation is driven by motivation but refined through inhibitory signals from neurons associated with competing or irrelevant decision options. Critically, the spontaneous variability inherent in this process affords the potential for the exploration of novel patterns, and through neuroplasticity, the potential to write and rewrite the synaptic structures that channel neuro-electrical activity in the first place. Seen in this light, our neural network is a “written artifact”, representing the history of useful sculptings of an endless ocean of spontaneous activity – the manifestation of evolutionary, developmental, and experiential unfolding. This written artifact is similar to what German philosopher Friedrich Schelling called architecture as “frozen music” – in our example, the music is the ongoing spontaneous activity itself and the synaptic strengths the resultant architecture (Schelling, 1845).
@@ 4  The Medium Shapes the Message
This discussion highlights our emphasis on the word medium: the defining mediation between subject and object. A medium’s spatiotemporal features are key in this mediation process. Canadian philosopher Marshall McLuhan famously coined the phrase, “The medium is the message”, observing that the unique properties associated with a novel medium profoundly shape our conscious experience – they literally become stretched “extensions” of our perceptual process, as he put it. Consequently, the real message is not a code in a channel of information, but rather how a medium’s unique characteristics directly shape our experience from the ground up. For this reason, McLuhan later adopted the title “the medium is the massage”, to imply how the unique features of mediums “massage” our conscious experience (McLuhan, 1967).